OpenStudy (anonymous):

how to put -5 and 4+3i as zeros? Ex. -3 is a zero. (x+3) Ex. 5i is a zero. (x+5i)(x-5i)

OpenStudy (aum):

Both in one equation or two separate equations?

OpenStudy (anonymous):

3 separate

OpenStudy (aum):

But you have only two zeros. How three separate equations?

OpenStudy (anonymous):

i comes in pairs

OpenStudy (aum):

But that is not what I am asking. Are the given zeros to the same equation or separate equations.

OpenStudy (anonymous):

given to the same polynomial

OpenStudy (aum):

Exactly that was my original question. Proceed as follows:

OpenStudy (anonymous):

n=3 -5 and 4+3i are zeroes f(2)=91

OpenStudy (anonymous):

find the nth-degree polynomial function with real coefficients satisfying the given conditions

OpenStudy (aum):

-5 and 4+3i are zeros x = -5 x + 5 = 0. Therefore (x+5) is one factor. x = 4 + 3i x - 4 = 3i square both sides to get rid of the "i"" (x-4)^2 = (3i)^2 = -9 x^2 - 8x + 16 = 9 x^2 - 8x + 7 = 0. Therefore, (x^2 - 8x + 7) is another factor. Multiplying them together and introduce a constant in front of the polynomial for the more general case: The polynomial f(x) = a(x + 5)(x^2 - 8x + 7) f(2)=91 put x = 2 and f(x) = 91: 91 = a(2+5)(2^2 -8*2 + 7) 91 = a(7)(4-16+7) 91 = a(7)(-5) a = -91/35 f(x) =-91/35(x + 5)(x^2 - 8x + 7) multiply the (x + 5)(x^2 - 8x + 7) part. just double check my arithmetic to see if they are correct.

OpenStudy (anonymous):

thank you that helped a lot

OpenStudy (aum):

You are welcome.

OpenStudy (aum):

Just noticed a mistake above. Correcting it here: -5 and 4+3i are zeros x = -5 x + 5 = 0. Therefore (x+5) is one factor. x = 4 + 3i x - 4 = 3i Square both sides to get rid of the "i" (x-4)^2 = (3i)^2 = -9 x^2 - 8x + 16 = -9 x^2 - 8x + 25 = 0. Therefore, (x^2 - 8x + 25) is another factor. Multiplying the two factors together and introduce a constant in front of the polynomial for the more general case: The polynomial f(x) = a(x + 5)(x^2 - 8x + 25) f(2)=91 put x = 2 and f(x) = 91: 91 = a(2+5)(2^2 -8*2 + 25) 91 = a(7)(4-16+25) 91 = a(7)(13) 91 = 91a a = 1 f(x) = (x + 5)(x^2 - 8x + 25) f(x) = x^3 - 8x^2 + 25x + 5x^2 - 40x + 125 f(x) = x^3 - 3x^2 - 15x + 125 is the polynomial.